OpenStudy (anonymous):

solve for v v^2+2v-15=0

OpenStudy (anonymous):

(v+5)(v-3)=0 v+5=0 v=-5 v-3=0 v=3

OpenStudy (anonymous):

v = -5 or v = 3

hero (hero):

oops

OpenStudy (anonymous):

v^2+5v-3v-15=0 v(v+5)-3(v+5)=0 (v+5)(v-3)=0 v=-5,3

OpenStudy (anonymous):

"v^2+5v-3v-15=0" That's a very interesting way of looking at it... yet it works. I like it.